.

May 23 2025 Eric Babson babson@math.ucdavis.edu 1094 Cragmont Avenue Berkeley CA, 94708

Dear Mr President:

I am concerned about genocide.

You bear significant responsibility for the starvation in Gaza and are morally obliged to act. You have provided more financial, military and political support for Prime Minister Netanyahu than any other entity with billions in military aid expedited and a call to expel Palestinians just this year. Last month many congressional Democrats have finally called out the horror of the food blockade and the cynical plan to use aid for military ends. The plan continues to have your administration’s support while Republicans continue their silence. This week the aid plan was rolled out forcing thousands to travel long distances amid the destroyed infrastructure only to find little food and a slaughter by Israeli troops with accounts also putting American contractors on the scene as well.

Please assure me that you will use all means available to end the blockade of Gaza.

My pride in the US derives in part from our humanitarian operations.

Sincerely, Eric Babson